# Federated nnU-Net for Privacy-Preserving Medical Image Segmentation

### v1

@@ Abstract
The nnU-Net framework has played a crucial role in medical image segmentation and has become the gold standard in multitudes of applications targeting different diseases, organs, and modalities. However, so far it has been used primarily in a centralized approach where the data collected from hospitals are stored in one center and used to train the nnU-Net. This centralized approach has various limitations, such as leakage of sensitive patient information and violation of patient privacy. Federated learning is one of the approaches to train a segmentation model in a decentralized manner that helps preserve patient privacy. In this paper, we propose FednnU-Net, a federated learning extension of nnU-Net. We introduce two novel federated learning methods to the nnU-Net framework – Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg) – and experimentally show their consistent performance for breast, cardiac and fetal segmentation using 6 datasets representing samples from 18 institutions. Additionally, to further promote research and deployment of decentralized training in privacy-constrained institutions, we make our plug-n-play framework public. The source-code is available at https://github.com/faildeny/FednnUNet.
@@
1.  Introduction
Medical image segmentation plays a key role in clinical workflows to accurately delineate anatomical structures and help with diagnosis, treatment planning, and disease monitoring [1, 2, 3, 4, 5, 6]. Deep learning methods, especially convolutional neural networks (CNNs), have shown great success in this field. One of the most effective approaches is the no-new-U-Net framework (nnU-Net) [7] which, over the years, has established itself as the go-to framework for medical image segmentation task, evidenced by its multiple first-place finishes on public leaderboards and its widespread adoption as the foundation for further model development across diverse segmentation challenges [8, 9]. nnU-Net extends the widely adopted U-Net architecture [10], which utilizes an encoder-decoder structure with skip connections to capture both semantic and localization information. Going beyond the original U-Net, nnU-Net automatically configures network architectures tailored to a given dataset, provides a robust training schedule and data augmentation strategy, as well as postprocessing methods to further improve the final performance. This “no-new” principle eliminates much of the manual effort associated with hyperparameter tuning and architectural engineering, rendering nnU-Net both powerful and highly convenient across a broad range of biomedical tasks [7, 11, 12, 13]. Despite these advancements, the development of robust segmentation models often requires aggregating extensive datasets from multiple institutions to capture diverse patient populations and imaging protocols. However, the consolidation of such data is frequently impeded by stringent data privacy regulations, ethical considerations, and logistical challenges associated with data sharing. Federated learning (FL) [14] has emerged as a promising solution to these challenges by facilitating collaborative model training across multiple institutions without the need to exchange raw data. Federated learning, however, presents significant challenges when data are heterogeneous across institutions. Traditional FL approaches typically assume that all participants train identical models on similarly distributed data, yet in real-world scenarios, clients often have different computational resources and dataset characteristics. To address this disparity, methods such as HeteroFL [15] have been introduced to enable federated training among clients with varying model capacities, while more recent efforts [16] have focused on mitigating label noise and inconsistencies inherent in federated settings. Collectively, these approaches underscore the growing need for robust FL techniques that can simultaneously manage model and data heterogeneity. Regardless of the significant federated developments in medical domain [17, 18, 19, 20, 21, 22, 23], those methods have not been explicitly tailored for medical image segmentation tasks using nnU-Net. Moreover, nnU-Net’s plug-and-play design, which has made it the gold standard in centralized settings, no longer functions seamlessly in federated environments, necessitating tailored adaptations to overcome the challenges of decentralized, heterogeneous data. To bridge the gap between the decentralized learning paradigms and the robust capabilities of nnU-Net framework, we introduce FednnU-Net, federated extension of nnU-Net that enables collaborative training while preserving data privacy. To the best of our knowledge, it is the first fully-federated and privacy-preserving implementation of nnU-net. We propose two novel methods–Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg). These methods address the challenges posed by variations in dataset properties across different centers, ensuring that each institution maximizes the benefit from shared knowledge training. Our experiments demonstrate that both methods consistently outperform local, non-collaborative training setups for 2D and 3D segmentation tasks across six public and non-public datasets. By extending nnU-Net’s automated pipeline to federated learning and introducing methods to handle data heterogeneity, this work aims to advance the development of robust, scalable, and privacy-preserving medical image segmentation solutions.
2.  Methods
@@
We propose two novel methods to federate nnU-Net’s pipeline for image segmentation: FFE and AsymFedAvg. These two approaches address the challenges of decentralized, heterogeneous data in model training while maintaining nnU-Net’s self-configuring advantages.
2.1.  nnU-Net vs FednnU-Net
@@
nnU-Net is a multi-stage segmentation pipeline that begins with a preparation stage, where dataset fingerprint extraction is performed and a customized training plan (including a tailored network architecture) is generated. This is followed by a training stage and concludes with a post-training stage that selects the best-performing models and optimal image post-processing techniques. To extend nnU-Net into a federated learning setting, we integrate our proposed methods at key points in this pipeline. The first method, FFE, is incorporated into the preparation stage to ensure that data characteristics from all participating institutions are captured and considered when generating the training plan. The second method, AsymFedAvg, operates during the training stage and facilitates the aggregation of model weight updates across nodes with heterogeneous architectures. The following sections provide detailed descriptions of each method. For clarity, we use the terms ”node” and ”client” interchangeably to refer to each institution participating in the decentralized training process, while ”server” describes the coordinating agent responsible for communication and aggregation of inputs from the nodes.
2.2.  Federated Fingerprint Extraction
@@
The proposed FFE method enables the analysis of the federated datasets and the configuration of nnU-Net, allowing to determine a unified, one-fits-all training strategy for all data centers in the federated setup. This approach effectively approximates the configuration behaviour in a centralized environment. The dataset’s fingerprint is a crucial component of the nnU-Net’s adaptive configuration, summarizing the key dataset characteristics that are used to determine the model configuration and preprocessing strategies. The FFE approach comprises two key steps:
  • Local Fingerprint Extraction: Each participating node generates their dataset fingerprint by invoking the extraction process of nnU-Net. This local fingerprint encloses the local dataset’s pixel intensity-based metrics, as well as spatial characteristics (such as spacing and voxel size) (see Table 1).
  • Global Fingerprint Aggregation:
    The nodes share their local fingerprint with the server, where these get aggregated to form a global fingerprint. Explicitly, the characteristics are aggregated as follows:
    $\hat{F}={\hat{f}k=g_{k}({f_{k}^{i}}{i=1}^{N})|k\in\mathcal{K}}$ (1)
    where $\hat{F}$ denotes the aggregated fingerprint, $f_{k}^{i}$ represents the $k-th$ fingerprint key from node $i$, and $g_{k}$ is the corresponding aggregation function. Table 1 lists the corresponding $g_{k}$ functions for each fingerprint key. The weighted_mean is calculated with weights proportional to data samples quantity from each participating node.
Table 1: Dataset fingerprint characteristics and aggregation Functions.
Fingerprint Key Aggregation Function $g_{k}$
Pixel intensity
max global max
min global min
mean weighted_mean
median weighted_mean
std weighted_mean
percentile_00_5 weighted_mean
percentile_99_5 weighted_mean
Shape
median_relative_size_after_cropping weighted_mean
shapes_after_crop concatenate
spacings concatenate
Fig. 1 showcases an overview of the FFE process pipeline. As a final step, each node calculates its own nnU-Net training plan based on the received global fingerprint and local hardware characteristics. This process ensures that all nodes will be optimizing either highly compatible or, in most cases, identical network architectures (if the default nnU-Net’s memory constraint has not been altered between nodes).
2.3.  Asymmetric Federated Averaging
In scenarios where nnU-Net is deployed across federated nodes, variations in model architecture can arise due to differences in data characteristics or computational capabilities. To address this inherent heterogeneity, we extend the classical FedAvg approach by developing an Asymmetric Federated Averaging strategy (AsymFedAvg). This method enables nodes to share only the parts of the model that have a common space with the rest of the federation, effectively eliminating the need to constrain all nodes to a single, identical architecture.
At each participating node, we initialize nnU-Net’s adaptive configuration pipeline to generate node-specific U-Net architectures based on the locally extracted dataset fingerprints of the centers. A set of example configurations obtained for each data center is presented in Table 2.
@@ 2.3.  Asymmetric Federated Averaging
$\small\mathcal{C}({S^{k}}_{k=1}^{K})={l|\forall k,j\in\mathcal{K}:l\in\mathcal{L}^{k}\cap\mathcal{L}^{j}\land\text{dim}(\theta_{l}^{k})=\text{dim}(\theta_{l}^{j})}$ (3)
@@
$\hat{\theta}_{l}=\frac{1}{|\mathcal{K}l|}\sum{k\in\mathcal{K}_{l}}\theta_{l}^{k}$ (4)
@@
$\hat{S}={(l,\hat{\theta}l)|l\in\mathcal{C}({S^{k}}{k=1}^{K})}$ (5)
During each federation round $t$, the update process goes as follows:
@@
The AsymFedAvg allows nodes to maintain their unique architectural components while sharing knowledge via their common space (compatible layers), handling the diverse nature of medical imaging data across different centers. Fig. 2 depicts an overview of the AsymFedAvg pipeline and the process of updating the heterogeneous architectures in each training round.
Figure 2: The Asymmetric Federated Averaging (AsymFedAvg) Pipeline process. The nodes of the federated setup generate a local nnU-Net from the nnU-Net pipeline. During training, the server receives the local nnU-Net weights of each node and aggregates them utilizing the Aggregation Function of AsymFedAvg. Finaly, at the end of each training round, the server distributes the aggregated layers of the network back into the nodes. This process happens throughout the whole training.
2.4.  Implementation
2.4.1 Federated Learning Environment
For the federated communication, we select the Flower framework [24], which is an open-source library designed to streamline federated learning tasks on a cluster of machines. Flower’s high-level API eliminates the need for delving into complex and time-consuming processes such as network communication between nodes and model aggregation, as it already supports multiple FL optimization algorithms such as Federated Averaging (FedAvg), FedProx [25] and Federated SGD (FedSGD) [14].
Additionally, Flower’s client-server architecture is well-suited for our multi-center setup, as it allows seamless integration of local models while maintaining centralized coordination on the server.
2.4.2 FednnU-Net
Our framework maintains full functionality of the original nnU-Net while offering a federated extension that allows experimentation in both real and simulated environments. Its modular structure enables seamless integration not only with the original nnU-Net but also with its customized versions, allowing researchers to adapt their traditional centralized training methods to a federated setup.
2.4.3 Distributed Setup
@@
Node CPU (Cores/GHz)* GPU (Memory)** OS ***
Node 1 i7-8700K (12/3.20) GTX 1080 (8GB) Ubuntu 22.04
Node 2 i7-8700K (12/3.20) GTX 1080 (8GB) Windows 10
Node 3 i7-9700K (8/3.60) RTX 2080S (8GB) Ubuntu 22.04
Node 4 i7-9700 (8/3.00) RTX 2080S (8GB) Ubuntu 22.04
*GPU: Graphics Processing Unit, **CPU: Central Processing Unit, ***OS: Operating system
3.  Datasets
We evaluate our proposed methods on three distinct clinical applications, using datasets that serve as valuable testbeds due to their multi-institutional and heterogeneous cohorts. The following sections provide a detailed description of the datasets employed in each use case.
(a) MAMA-MIA
(b) M&M
(c) AIMIX
Figure 3: Example images with annotations from each data center used in the study.
3.1.  Breast MRI
3.1.1 MAMA-MIA
The MAMA-MIA dataset [26] is the largest publicly available collection of breast dynamic contrast-enhanced MRI (DCE-MRI) cases with expert primary tumor segmentations, providing an essential resource for advancing and benchmarking AI algorithms in breast cancer imaging. It comprises 1,506 cases acquired between 1995 and 2016 from four cohorts within The Cancer Imaging Archive (TCIA): ISPY-1 (n=171), ISPY-2 (n=980), Duke (n=291), and NACT (n=64). The annotation process focused on delineating primary tumors and non-mass-enhanced lesions within a single breast, using the first post-contrast phase as the reference—a phase typically associated with maximum enhancement of cancerous tissues that offers superior contrast between lesions and surrounding healthy tissue compared to later phases. However, significant imaging differences among these cohorts (Table 4) pose challenges for training robust models. For example, the NACT cohort predominantly features scans with a 2 mm slice thickness and an average of 60 slices per case, whereas Duke’s samples are mostly acquired with 1 mm thickness and include an average of 169 slices per case. Moreover, the centers used different scanner models and field strengths, further contributing to the variability in imaging protocols.
@@ 3.1.1 MAMA-MIA
3.1.2 EuCanImage:
Through the European project EuCanImage [^1] https://eucanimage.eu/, we access a non-public breast MRI collection comprising 574 samples from three different countries. The EuCanImage dataset represents a multi-institutional effort to create standardized imaging biobanks for cancer research and serves here as an external test set to validate our methodology on data acquired under protocols that differ markedly from those in MAMA-MIA. Detailed specifications of the acquisition parameters are provided in Table 5. Notably, while MAMA-MIA includes both axial and sagittal acquisitions, the EuCanImage collection consists exclusively of axial slices. In addition, images in EuCanImage are predominantly acquired from scanner models that differ from those represented in MAMA-MIA, and a significantly larger fraction of the scans are captured at 3T field strength. Although the spatial image characteristics are largely similar between the two datasets, these differences in scanner models, field strengths, and imaging orientations lead to distinct scan appearances, potentially challenging the robustness of segmentation algorithms.
@@ 3.1.2 EuCanImage:
3.2.  Cardiac MRI
3.2.1 M&Ms
To further validate our approach, we leverage the Multi-center, Multi-vendor & Multi-disease (M&Ms) cardiac MRI datasets. As a part of our study, we curate a merged version of datasets that were used in the M&Ms challenges [27];[28] by excluding the overlapping samples based on anonymized patient IDs provided by each institution. This curation allows to reach the full potential of this data collection without the risk of replicated samples. The final, accumulated set consists of 543 unique, short-axis cardiac MRI scans with expert annotations from 5 distinct institutions. We make this collection publicly available [^2] This data resource will be released upon the acceptance of this article.. The M&Ms datasets are particularly valuable for validation due to their inherent heterogeneity (Table 6), they include images from four different vendors (Siemens, GE, Philips, Canon), various scanner models (from older systems like Symphony to newer ones like Vantage Orian), and different field strengths (1.5T and 3.0T). This multi-center, multi-vendor nature mirrors the real-world variability in clinical settings, making these datasets ideal for testing the generalizability of our approach.
@@ 3.2.1 M&Ms
Hospital Vall d’Hebron Siemens Symphony TIM (63.18%) Avanto Fit (15.48%) Other (21.34%) 9.5 13 1.5 (97.91%) 3.0 (2.09%) 239
Clinica Sagrada Familia Philips Achieva 10 13 1.5 98
Universitätsklinikum Hamburg-Eppendorf Philips Achieva 10 11 1.5 51
Hospital Universitari Dexeus GE Signa Excite (75.24%) Signa HDxt (23.81%) Signa Explorer (0.95%) 9.9 12 1.5 (88.57%) 3.0 (11.43%) 105
@@ 3.2.2 ACDC:
3.3.  Fetal Ultrasound
3.3.1 AIMIX
With a focus on resource-constrained environments, the AIMIX [^3] https://aimix-erc.eu/ dataset was developed to advance inclusive image analysis in prenatal care. This dataset addresses critical challenges, including the scarcity of African medical data, limited computational resources, and a lack of clinical expertise. The dataset comprises pregnancy blind-sweeps 2D ultrasound scans, along with associated patient metadata such as age and ethnicity. The data span multiple international sites, including Algeria, Egypt, Malawi and Spain, reflecting diverse geographic and healthcare settings.
In this study, we focus on abdominal segmentation, a crucial step in prenatal ultrasound analysis. To enable the evaluation on this multi-center dataset, we provide a set of 130 standard-plane images with masks corresponding to the fetal abdomen [^4] This data resource will be released upon the acceptance of this article.. Our data processing workflow involved detecting the optimal abdominal plane from the each blind-sweep and segmenting it using a 4-point ellipse-fitting algorithm [29], utilizing a semi-automatic software by a researcher with more than five years of experience and reviewed by an expert clinician. Table 7 lists the number of samples and the mean resolution for each center.
@@ 3.3.1 AIMIX
3.3.2 ACOUSLIC-AI Sierra Leone:

### v2

@@ Abstract
The nnU-Net framework has played a crucial role in medical image segmentation and has become the gold standard in multitudes of applications targeting different diseases, organs, and modalities. However, so far it has been used primarily in a centralized approach where the collected data is stored in the same location where nnU-Net is trained. This centralized approach has various limitations, such as potential leakage of sensitive patient information and violation of patient privacy. Federated learning has emerged as a key approach for training segmentation models in a decentralized manner, enabling collaborative development while prioritising patient privacy. In this paper, we propose FednnU-Net, a plug-and-play, federated learning extension of the nnU-Net framework. To this end, we contribute two federated methodologies to unlock decentralized training of nnU-Net, namely, Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg). We conduct a comprehensive set of experiments demonstrating high and consistent performance of our methods for breast, cardiac and fetal segmentation based on a multi-modal collection of 6 datasets representing samples from 18 different institutions. To democratize research as well as real-world deployments of decentralized training in clinical centres, we publicly share our framework at https://github.com/faildeny/FednnUNet.
@@
Introduction
Medical image segmentation plays a key role in clinical workflows to accurately delineate anatomical structures and help with diagnosis, treatment planning, and disease monitoring [1, 2, 3, 4, 5]. Deep learning methods, especially convolutional neural networks (CNNs) [6, 7], have shown great success in this field. One of the most effective approaches is the no-new-U-Net framework (nnU-Net) [8] which, over the years, has established itself as the go-to framework for medical image segmentation task, evidenced by its multiple first-place finishes on public leaderboards and its widespread adoption as the foundation for further model development across diverse segmentation challenges [9, 10].
nnU-Net extends the widely adopted U-Net architecture [11], which utilizes an encoder-decoder structure with skip connections to capture both semantic and localization information. Going beyond the original U-Net, nnU-Net automatically configures network architectures tailored to a given dataset, provides a robust training schedule and data augmentation strategy, as well as postprocessing methods to further improve the final performance. This “no-new” (nn) principle largely eliminates the manual effort associated with hyperparameter tuning and architectural engineering, rendering nnU-Net both powerful and highly convenient across a broad range of biomedical tasks [8, 12, 13, 14].
Despite these advancements, the development of robust segmentation models often requires aggregating extensive datasets from multiple institutions to capture diverse patient populations and imaging protocols. However, the consolidation of such data is frequently impeded by stringent data privacy regulations, ethical considerations, and logistical challenges associated with data sharing. Federated learning (FL) [15] was introduced as a promising solution to these challenges by facilitating collaborative model training across multiple institutions without the need to exchange raw data.
Although nnU-Net has become a standard for medical image segmentation in centralized settings, its use in FL is challenging due to its dataset-specific customization which involves preprocessing, normalization, model choice, and training schedules based on the underlying data [8]. The model architecture itself includes design decisions regarding the size and stride of convolutional kernels, as well as the number of network blocks. Such level of customization fundamentally conflicts with standard FL methods like FedAvg [15] that require uniform model architectures and training procedures across clients.
To solve these limitations, we introduce FednnU-Net–the first fully federated privacy-preserving implementation of nnU-Net. FednnU-Net is specifically designed to maintain the automated and out-of-the-box nature of nnU-Net while overcoming the challenges associated with decentralized, heterogeneous data.
To this end, we propose two methods—Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg) to unlock distributed training of nnU-Net. These methods address the challenges posed by variations in dataset properties across different centers, ensuring that each institution maximizes the benefit from shared knowledge training. Our experiments demonstrate that both methods consistently outperform local, non-collaborative training setups for 2D and 3D segmentation tasks across six public and non-public datasets.
By extending nnU-Net’s automated pipeline to FL and introducing methods to handle data heterogeneity, this work advances the development of robust, scalable, and privacy-preserving medical image segmentation solutions. Furthermore, its modular design also unlocks access to bringing a wide range of nnU-Net extensions to the federated settings, including improved training efficiency through e.g. curriculum learning [16], enhanced model explainability [17], continual learning [18], or uncertainty estimation [19, 20, 21].
Related Work
Federated learning in medical imaging
FL has been widely adopted in medical imaging for tasks such as disease classification, lesion detection, and organ segmentation across multiple institutions [22, 23, 24]. A range of federated segmentation methods have been introduced to tackle key challenges such as data heterogeneity, communication inefficiency, and limited generalization across sites. Notable examples include FedEvi [25], which employs evidence-based aggregation to account for uncertainty in model updates; FedGS [26], which utilizes gradient statistics to enhance robustness; and Fed-MENU [27], which integrates model ensembling with uncertainty-aware learning to improve stability. Transformer-based [28] methods leverage self-attention mechanisms and present high performance in distributed training for positron emission tomography (PET) and computed tomography (CT) segmentation [29]. Additionally, recent approaches like FedFMS [30] incorporate foundation models like SAM [31] and MedSAM [32] into FL through fine-tuning, aiming to harness large-scale pretrained representations for downstream segmentation tasks in a privacy-preserving context.
Despite the overall progress, these methods present limitations in terms of practical applicability. Many rely on highly customized architectures, complex aggregation strategies, or specialized pretraining, which can pose significant barriers for researchers without profound expertise in machine learning. In contrast to nnU-Net framework, Transformer- and foundational model-based approaches offer much less flexibility in terms of computational resources, making them less feasible in typical clinical environments with constrained hardware.
These constraints limit their out-of-the-box usability and, as a result, chances of clinical application. There remains a strong need for universal FL methods that are robust across datasets, computationally efficient, and readily applicable by a broad range of users—including those with limited resources or domain-specific requirements.
Heterogeneous model aggregation
Conventional FL assumes identical models and similarly distributed data across clients. However, real-world deployments often involve heterogeneous datasets and varying compute capacities. To address this, HeteroFL [33] allows training subnetworks of different widths from a shared supernetwork. While effective for standard CNNs, it is incompatible with nnU-Net, which adapts model depth, resolution, and preprocessing to dataset-specific features. FjORD [34] introduces Ordered Dropout for training nested submodels, enabling dynamic scaling. Yet, it relies on a fixed architecture, limiting its applicability to nnU-Net’s dataset-driven design. FedRolex [35] improves fairness by cyclically training submodels from a global network. However, it still assumes all client models are subsets of a shared template—an assumption that nnU-Net’s self-configuring framework violates. Since none of the existing methods accommodates the flexible and adaptive nature of nnU-Net, we develop a new federated approach tailored specifically to nnU-Net’s architecture and requirements.
Methods
@@
We propose two tailored methods to federate nnU-Net’s pipeline for image segmentation, namely, FFE and AsymFedAvg. These two approaches are tailored towards addressing the challenges of decentralized, heterogeneous data in model training while simultaneously allowing to maintain nnU-Net’s self-configuring advantages.
nnU-Net vs FednnU-Net
@@
nnU-Net is a multi-stage segmentation pipeline initiated by a preparation stage, where dataset fingerprint extraction is performed and a customized training plan is generated, which includes a tailored segmentation network architecture. This is followed by a training stage before concluding with a post-training stage that selects the best-performing models alongside optimal image post-processing techniques. To extend nnU-Net into a FL framework, we integrate our proposed methods in the respectively corresponding components of this pipeline. The first method, FFE, is incorporated into the preparation stage, ensuring that data characteristics from all participating institutions are captured and considered when generating the training plan. The second method, AsymFedAvg, operates during the training stage and facilitates the aggregation of model weight updates across nodes with heterogeneous network architectures. With the following sections providing detailed descriptions of each method, we use the terms ”node” and ”client” interchangeably to refer to each institution participating in the decentralized training process. We further denote the term ”server” as the coordinating agent responsible for communication and aggregation of inputs from the nodes.
Federated Fingerprint Extraction
The proposed FFE method enables the analysis of the federated datasets and the configuration of nnU-Net, allowing to determine a unified, one-fits-all training strategy for all data centers in the federated setup. This approach effectively approximates the configuration behavior in a centralized environment. The dataset’s fingerprint is a crucial component of the nnU-Net’s adaptive configuration, summarizing the key dataset characteristics that are used to determine the model configuration and preprocessing strategies. The FFE approach comprises two key steps:
  • Local Fingerprint Extraction: Each participating node generates their dataset fingerprint by invoking the extraction process of nnU-Net. This local fingerprint encloses the local dataset’s spatial characteristics such as spacing and voxel size.
  • Global Fingerprint Aggregation: The nodes share their local fingerprint with the server, where these get aggregated to form a global federated fingerprint z. Specifically, new lists of shapes_after_crop and spacings are created by the concatenation of their local representations.
@@
Fig. 1 showcases an overview of the FFE process pipeline. As a final step, each node calculates its own nnU-Net training plan based on the combination of the received global fingerprint and the local hardware characteristics. This process ensures that all nodes will be optimizing either highly compatible or, in the case of nodes with sufficiently similar local memory and hardware constraints, identical network architectures.
Asymmetric Federated Averaging
In scenarios where nnU-Net is deployed across federated nodes, variations in model architecture can arise due to differences in data characteristics or computational capabilities. To address this inherent heterogeneity, we extend the classical FedAvg approach by developing a novel Asymmetric Federated Averaging strategy (AsymFedAvg). In contrast to other heterogeneous aggregation methods mentioned in Heterogeneous model aggregation, our method allows the aggregation of models with different lengths and layer configurations. AsymFedAvg enables nodes to share only the parts of the model that are common across the rest of the federation, effectively eliminating the need to constrain all nodes to a single, identical architecture.
At each participating node, we initialize nnU-Net’s adaptive configuration pipeline to generate node-specific U-Net architectures based on the aforementioned locally extracted dataset fingerprints of the centers. For illustration, a set of example configurations obtained from each data center is presented in Table 1.
Figure 2: Overview of the Asymmetric Federated Averaging (AsymFedAvg) Training Pipeline. The nodes of the federated setup generate a local nnU-Net from the nnU-Net pipeline. During training, the server receives the local nnU-Net weights of each node and aggregates them utilizing the Aggregation Function of AsymFedAvg. At the end of each training round, the server distributes the aggregated layers of the network back into the nodes. This process is respectively repeated throughout the entire model training process.
@@
$\mathcal{C}({S^{k}}_{k=1}^{K})={l|\forall k,j\in\mathcal{K}:l\in\mathcal{L}^{k}\cap\mathcal{L}^{j}\land\text{dim}(\theta_{l}^{k})=\text{dim}(\theta_{l}^{j})}.$ (2)
@@
$\hat{\theta}_{l}=\frac{1}{|\mathcal{K}l|}\sum{k\in\mathcal{K}_{l}}\theta_{l}^{k},$ (3)
@@
$\hat{S}={(l,\hat{\theta}l)|l\in\mathcal{C}({S^{k}}{k=1}^{K})}.$ (4)
During each federation round $t$, the update process goes as presented in Algorithm  1.
@@
Notation:
  • $S^{k}_{t}$: State dictionary of node $k$ at round $t$
  • $\mathcal{C}$: Layer compatibility function
  • $\hat{\theta}_{l,t}$: Aggregated parameters for layer $l$
The AsymFedAvg allows nodes to maintain their unique architectural components while sharing knowledge via their common space of compatible layers, handling the diverse nature of medical imaging data across different centers. Fig. 2 depicts an overview of the AsymFedAvg pipeline and illustrates the process of updating the heterogeneous architectures in each training round.
Implementation
Federated Learning Environment
For the federated communication, we select the Flower framework [36], which is an open-source library designed to streamline FL tasks on a cluster of machines. Flower’s high-level application programming interface (API) eliminates the need for delving into complex and time-consuming processes such as network communication between nodes and model aggregation, as it supports multiple FL optimization algorithms such as Federated Averaging (FedAvg), FedProx [37] and Federated SGD (FedSGD) [15]. Additionally, Flower’s client-server architecture is well-suited for our multi-center setup, facilitating seamless integration of local models while maintaining centralized coordination on the server.
FednnU-Net Framework
Our framework maintains full functionality of the original nnU-Net while offering a federated extension that allows experimentation in both real-world and simulated environments. Focusing on usability, its modular structure enables seamless integration not only with the original nnU-Net but also with respective customized versions, allowing researchers to adapt their traditional centralized training methods to a federated setup. Key features include:
  • Custom Architecture Selection: Users can select from nnUNet’s built-in architecture presets, including configurations optimized for high performance (e.g., residual encoder variants), even if they require more memory or compute.
  • Per-Client Resource Configuration: GPU memory constraints can be individually defined for each client, enabling heterogeneous resource allocation scenarios such as in AsymFedAvg.
  • Training Hyperparameter Control: Users can configure learning rate schedules, loss functions, optimizer settings, and number of training epochs through the native nnUNetTrainer class.
  • Support for Pretrained Weights: Initialization from pretrained weights is supported, allowing faster convergence or domain adaptation across clients.
  • Simulated and Real-World Federation: The same codebase supports both simulated federated setups (single machine, multi-client emulation) and deployment across physically distributed clients.
This design ensures both high flexibility for research experimentation and practicality for real-world deployment across diverse clinical environments.
Distributed Setup
@@
Node CPU (Cores/GHz) GPU (Memory) OS§
Node 1 Intel i7-8700K (12/3.20) NVIDIA GTX 1080 (8GB) Ubuntu 22.04
Node 2 Intel i7-8700K (12/3.20) NVIDIA GTX 1080 (8GB) Windows 10
Node 3 Intel i7-9700K (8/3.60) NVIDIA RTX 2080S (8GB) Ubuntu 22.04
Node 4 Intel i7-9700 (8/3.00) NVIDIA RTX 2080S (8GB) Ubuntu 22.04
CPU: Central Processing Unit.  GPU: Graphics Processing Unit.  §OS: Operating System.
Datasets
We evaluate our proposed methods on three distinct clinical applications (i.e., breast, cardiac, fetal), using datasets that serve as valuable testbeds due to their multi-institutional and heterogeneous cohorts. The following sections provide a detailed description of the datasets employed in each use case.
Figure 3: Example images with annotations from each data center used in this work. From the top: Breast MRI, Cardiac MRI, Fetal US
Breast MRI
MAMA-MIA
The MAMA-MIA dataset [38] is the largest publicly available collection of breast dynamic contrast-enhanced MRI (DCE-MRI) cases with expert primary tumor segmentations, providing an essential resource for advancing and benchmarking artificial intelligence (AI) algorithms in breast cancer imaging. It comprises 1,506 cases acquired between 1995 and 2016 from four cohorts within The Cancer Imaging Archive (TCIA): ISPY-1 (n=171), ISPY-2 (n=980), Duke (n=291), and NACT (n=64). The annotation process focused on delineating primary tumors and non-mass-enhanced lesions within a single breast, using the first post-contrast phase as the reference—a phase typically associated with maximum enhancement of cancerous tissues that offers superior contrast between lesions and surrounding healthy tissue compared to later phases. However, significant imaging differences among these cohorts (Table 3) pose challenges for training robust models. For example, the NACT cohort predominantly features scans with a 2 mm slice thickness and an average of 60 slices per case, whereas Duke’s samples are mostly acquired with 1 mm thickness and include an average of 169 slices per case. Moreover, the centers used different scanner models and field strengths, further contributing to the variability in imaging protocols.
@@
EuCanImage
Through the European project EuCanImage [39], we access a non-public breast MRI collection comprising 574 samples from three different countries. The EuCanImage dataset represents a multi-institutional effort to create standardized imaging biobanks for cancer research and serves here as an external test set to validate our methodology on data acquired under protocols that differ markedly from those in MAMA-MIA. Detailed specifications of the acquisition parameters are provided in Table 4. Notably, while MAMA-MIA includes both axial and sagittal acquisitions, the EuCanImage collection consists exclusively of axial slices. In addition, images in EuCanImage are predominantly acquired from scanner models that differ from those represented in MAMA-MIA, and a significantly larger fraction of the scans are captured at 3T field strength. Although the spatial image characteristics are largely similar between the two datasets, these differences in scanner models, field strengths, and imaging orientations lead to distinct scan appearances, potentially challenging the robustness of segmentation algorithms.
@@
Cardiac MRI
M&Ms
To further validate our approach, we leverage the Multi-center, Multi-vendor & Multi-disease (M&Ms) cardiac MRI datasets. As a part of our study, we curate a merged version of datasets that were used in the M&Ms challenges [40, 41] by excluding the overlapping samples based on anonymized patient IDs provided by each institution. This curation allows to reach the full potential of this data collection without the risk of replicated samples. The final, accumulated set consists of 543 unique, short-axis cardiac MRI scans with expert annotations from 5 distinct international institutions. We make this collection publicly available.
The M&Ms datasets are particularly valuable for validation due to their inherent heterogeneity (Table 5), they include images from four different vendors (Siemens, GE, Philips, Canon), various scanner models (from older systems like Symphony to newer ones like Vantage Orian), and different field strengths (1.5T and 3.0T). This multi-center, multi-vendor nature mirrors the real-world variability in clinical settings, making these datasets ideal for testing the generalizability of our approach.
@@
Hospital Vall d’Hebron Siemens
Symphony TIM (63.18%)
Avanto Fit (15.48%)
Other (21.34%)
9.5 13
1.5 (97.91%)
3.0 (2.09%)
239
Clinica Sagrada
Familia
Philips Achieva 10 13 1.5 98
Universitätsklinikum
Hamburg-Eppendorf
Philips Achieva 10 11 1.5 51
Hospital Universitari Dexeus GE
Signa Excite (75.24%)
Signa HDxt (23.81%)
Signa Explorer (0.95%)
9.9 12
1.5 (88.57%)
3.0 (11.43%)
105
@@
Fetal Ultrasound
AIMIX
With a focus on resource-constrained environments, the AIMIX [42] dataset was developed to advance inclusive image analysis in prenatal care. This dataset addresses critical challenges, including the scarcity of African medical data, limited computational resources, and a lack of clinical expertise. The dataset comprises pregnancy blind-sweeps 2D ultrasound scans, along with associated patient metadata such as age and ethnicity. The data span multiple international sites, including Algeria, Egypt, Malawi and Spain, reflecting diverse geographic and healthcare settings.
In this study, we focus on abdominal segmentation, a crucial step in prenatal ultrasound analysis. To enable the evaluation on this multi-center dataset, we provide a set of 308 standard-plane images with masks corresponding to the fetal abdomen. Our data processing workflow involved detecting the optimal abdominal plane from each blind-sweep and segmenting it using a 4-point ellipse-fitting algorithm [43], utilizing a semi-automatic software by a researcher with more than five years of experience and reviewed by an expert clinician. Table 6 lists the number of samples and the mean resolution for each center.
@@
ACOUSLIC-AI Sierra Leone